OpenStudy (anonymous):

a= -2 b= -3 c= 2 x= 2.1 y= 3 and z= -4.2 -I 5a + c I + I 3y + 2z I

OpenStudy (anonymous):

-(8)+(.6)=-7.4

OpenStudy (anonymous):

and is that the answer? can i get it step by step?

OpenStudy (anonymous):

\[-\left| 5(-2)+2 \right| + \left| 3(3)+2(-4.2) \right|\]

OpenStudy (anonymous):

\[-\left| -8 \right|+\left| 9-8.4 \right|\]

OpenStudy (anonymous):

\[-8+.6=-7.4\]

OpenStudy (anonymous):

final answer -7.4